OpenStudy (anonymous):

find the volume of the solid obtained by rotating the region bounded by y=(x^2)/4 x=2 y=0 about the y-axis

OpenStudy (tkhunny):

The whole cyllindrical solid is \(\pi\cdot 2^{2}\cdot 1 = 4\pi\). What's the interior part that you must discard?

OpenStudy (anonymous):

just that was the easy part . . . I will try to explain where I need help my math professor was a little confusing, I just do not full understand how to set up the integral, I drew a picture and everything but just cannot put the pieces together. Once I have the integral I can calculate it fine

OpenStudy (anonymous):

so would it be \[\pi \int\limits_{0}^{1}2-2\sqrt{y} dy\]

OpenStudy (tkhunny):

Too bad this is not an evaluate-the-integral problem. It is a set-up-the-integral problem. Think on this formula: Volume of Right Circular Solid: A = \(\pi r^{2}h\) \(\dfrac{\partial A}{\partial r} = 2\pi rh\) \(\dfrac{\partial A}{\partial h} = \pi r^{2}\)

OpenStudy (anonymous):

This is why I find this teacher so confusing he does one very basic general diagram and then goes striaght into exampes . . . he never even mentioned the formula you are talking about, I think (he also has a very heavy Russian accent)

OpenStudy (anonymous):

can you tell me what specific part of my integral is wrong so that I know where to look?

OpenStudy (tkhunny):

No one tends to mention that relatioship. It's there, though.

OpenStudy (tkhunny):

Why do you think there is anything wrong with your integral? Can you do it the other way?

OpenStudy (anonymous):

I just get confused when it is rotated about the y axis and sometimes mix up my limits of integration and the outside function also I didn't have the integral when I posted the question it just dawned on me when I posted it.

OpenStudy (tkhunny):

Well, since you are using Disks, there should be a single argument and it should be squared. Please refer to my second mysterious formula. Note: \(\pi\int\limits_{0}^{1}2\;dy = 2\pi\). That is NOT \(4\pi\) as would be correct for the entire cyllindrical solid. However, \(\pi\int\limits_{0}^{1}2^{2}\;dy = 4\pi\). Now THAT show promise. If you ponder my mysterious formulas, you will see it and you will not ever have to be confused.

OpenStudy (tkhunny):

Seriously, just square your arguments \(2^{2}\) and \((2\sqrt{y})^{2}\)